# Bootstrapped Control Limits for Score-Based Concept Drift Control Charts

### v1

@@ Abstract
Monitoring for changes in a predictive relationship represented by a fitted supervised learning model (aka concept drift detection) is a widespread problem, e.g., for retrospective analysis to determine whether the predictive relationship was stable over the training data, for prospective analysis to determine when it is time to update the predictive model, for quality control of processes whose behavior can be characterized by a predictive relationship, etc. A general and powerful Fisher score-based concept drift approach has recently been proposed, in which concept drift detection reduces to detecting changes in the mean of the model’s score vector using a multivariate exponentially weighted moving average (MEWMA). To implement the approach, the initial data must be split into two subsets. The first subset serves as the training sample to which the model is fit, and the second subset serves as an out-of-sample test set from which the MEWMA control limit (CL) is determined. In this paper we develop a novel bootstrap procedure for computing the CL. Our bootstrap CL provides much more accurate control of false-alarm rate, especially when the sample size and/or false-alarm rate is small. It also allows the entire initial sample to be used for training, resulting in a more accurate fitted supervised learning model. We show that a standard nested bootstrap (inner loop accounting for future data variability and outer loop accounting for training sample variability) substantially underestimates variability and develop a 632-like correction that appropriately accounts for this. We demonstrate the advantages with numerical examples.
Keywords: Control charts, Concept drift, Bootstrap, Predictive modeling, Machine learning.
@@ 1 Introduction
The increasing reliance on data-driven decision making has led to the widespread adoption of supervised learning models across various domains. These models aim to capture the predictive relationship $\mathbb{P}(Y|\mathbf{X})$ between a response variable $Y$ and covariates $\mathbf{X}$. However, a fundamental challenge arises when the predictive relationship in new data deviates from that used to train the model, potentially rendering the model’s predictions unreliable or obsolete [10, 18, 20], and/or reflecting a change in process behavior that should be detected. This challenge is particularly acute in domains like finance [16] and healthcare analytics [14], as most AI systems and algorithms require training data that may contain inherent biases or may not remain representative of the broader population over time [4].
@@
Statistical process monitoring (SPM) and control charts have long served as foundational tools for detecting changes in process characteristics over time [11]. The field has evolved considerably, with recent advances incorporating machine learning methods like kernel methods [2] and neural networks [13] to enhance monitoring capabilities. As datasets have grown larger and organizations increasingly rely on predictive models, the predictive relationship between variables has emerged as a critical characteristic that requires monitoring for quality control purposes, alongside traditional process variables monitoring.
@@
This evolution in the goals of monitoring reflects a fundamental shift in how data are used in practice and/or representing a change in process behavior that should be detected. While traditional SPM methods are designed to detect shifts in the distribution of process variables, modern applications often require understanding changes in the predictive relationships themselves. This phenomenon, known as concept drift in the machine learning literature, occurs when the relationship between input features $\mathbf{X}$ and the target variable $Y$ evolves over time [18], potentially degrading model performance [12, 18, 21] and/or representing a change in process behavior that should be detected. These changes can manifest gradually or abruptly, and may not affect predictive accuracy immediately [8], making them difficult to detect using conventional methods.
@@
Most existing concept drift detection methods fall into two categories, with each having significant limitations. The first category consists of error-based approaches that rely primarily on monitoring classification error rates or prediction accuracy metrics [3, 15]. While straightforward to implement, these methods often fail to detect concept drift when changes in $\mathbb{P}(Y|\mathbf{X})$ do not manifest as increased error rates, such as when decision boundaries shift in ways that maintain similar overall accuracy despite fundamental changes in the underlying relationship. The second category comprises adaptive learning algorithms, which continually retrain or update models to adapt to incoming data streams [8, 9, 17]. While these approaches can improve responsiveness, the adaptive model retraining also requires significant computation and may overfit transient changes.
@@
Related to, but distinctly different from concept drift detection, profile monitoring methods in statistical process control have been extensively developed to monitor functional relationships between response and predictor variables [19]. These methods typically involve monitoring regression parameters or fitted curves over time to detect changes in the relationship between a response variable and one or more explanatory variables [1, 6]. While this may appear similar to concept drift detection, profile monitoring fundamentally differs in its objectives and approaches and in the structure of the data to which it applies. In profile monitoring, data are grouped as batches of ($\mathbf{X}$, $Y$) observations, where some feature of $\mathbb{P}(Y|\mathbf{X})$ (typically $\mathbb{E}[Y|\mathbf{X}]$) for each batch represents a profile associated with the batch, and the objective is to monitor for changes in the nature of the profiles from batch to batch. This requires fitting predictive models separately to each batch of data. In contrast, in concept drift monitoring, data are individual ($\mathbf{X}$, $Y$) observations, and the objective is to detect whether the predictive relationship for new observations changes relative to what it was when some baseline model was fit to a prior set of training observations. Changes in the predictive relationship are detected by comparing a new stream of individual ($\mathbf{X}$, $Y$) observations to the baseline model, which does not involve fitting separate models to the new data.
@@
Recently, [20] introduced a new Fisher score-based concept drift detection that uses well-established statistical theory to show that detecting changes in $\mathbb{P}(Y|\mathbf{X})$ is equivalent to detecting changes in the mean of the score vector (the gradient of the log-likelihood) of the supervised learning model, a more familiar problem in SPM for which a conventional multivariate exponentially weighted moving average (MEWMA) control charts can be used. Their approach demonstrates superior detection power compared to traditional error-based concept drift methods, while also providing valuable diagnostic information to help understand the nature of detected changes. It also avoids having to continuously retrain the model, as in adaptive learning algorithms. Computing the MEWMA control limit (CL) in [20] requires that a second large sample of ($\mathbf{X}$, $Y$) observations be collected, in addition to the training sample to which the baseline model is fit. This can be achieved by splitting the training sample into two sets (with the baseline model fit to the first set, and the second set used to compute the CL), but this reduces the size of the data to which the baseline model is fit and generally requires very large sample sizes to accurately control false-alarm rate.
@@
Our main contribution is that we present a more data-efficient and reliable approach that uses bootstrapping to compute the CL from the same training sample to which the baseline model is fit, thereby allowing the entire sample to be used to train a more accurate baseline model, as well as providing much more accurate control of false-alarm rate. For this we develop a nested bootstrap procedure with the inner loop accounting for future data variability and outer loop accounting for training sample variability. This is challenging because a naïve nested bootstrap procedure substantially underestimates variability of the MEWMA monitoring statistic for reasons that we discuss later. To account for this, we develop a 632-like correction that controls false-alarm rate much more accurately than the two-sample CL calculation of [20], especially when the training sample size and/or the desired false-alarm rate is small. Our implementation takes advantage of the inherent parallelism in the nested bootstrap structure, making the method computationally efficient and practical for modern applications involving complex models like deep neural networks.
@@ 2 Background on Score-Based Concept Drift Detection
This work assumes a parametric supervised learning $g(\boldsymbol{\theta};\mathbf{X})$ is used to represent the conditional distribution $\mathbb{P}(Y|\mathbf{X};\boldsymbol{\theta})$ of the target variable $Y$ given the input features $\mathbf{X}\in\mathbb{R}^{p}$, where $\boldsymbol{\theta}$ denotes the parameters of the supervised learning model. For example, for classification problems, $g(\boldsymbol{\theta};\mathbf{X})$ directly outputs the response class probabilities; and for regression with Gaussian errors and squared-error loss, $g(\boldsymbol{\theta};\mathbf{X})$ represents the conditional mean $\mathbb{E}[Y|\mathbf{X};\boldsymbol{\theta}]$ of a Gaussian $\mathbb{P}(Y|\mathbf{X};\boldsymbol{\theta})$. As in [20], we assume $g(\boldsymbol{\theta};\mathbf{X})$ is fit to the training data via maximum likelihood estimation (MLE), perhaps with regularization.
Denote the training data to which $g(\boldsymbol{\theta};\mathbf{X})$ is fit by $\{(\mathbf{x}_{i},y_{i}):i=1,2,\cdots,n\}$, assumed to be an i.i.d. sample drawn from the joint distribution $\mathbb{P}(Y,\mathbf{X})$. For each observation $(\mathbf{x}_{i},y_{i})$, the (Fisher) score vector is defined as $\mathbf{s}(\boldsymbol{\theta};\mathbf{x}_{i},y_{i})=\nabla_{\boldsymbol{\theta}}\log\mathbb{P}(y_{i}|\mathbf{x}_{i};\boldsymbol{\theta})$. Under certain regularity conditions, if the assumed parametric model is indeed the true data-generating mechanism with true parameters denoted by $\boldsymbol{\theta}^{0}$, a fundamental result in statistical theory [5][Proposition 3.4.4] states that the expected value of the score vector, when evaluated at $\boldsymbol{\theta}^{0}$, is equal to zero
@@
In the context of machine learning, the notion of concept drift refers to the phenomenon where the underlying relationship between $\mathbf{X}$ and $Y$ evolves over time, which can be characterized as a shift in $\mathbb{P}(Y|\mathbf{X};\boldsymbol{\theta})$. In the parametric setting of [20], concept drift translates to a change in $\boldsymbol{\theta}$. Under certain identifiability conditions, when the parameters change (to some $\boldsymbol{\theta}\neq\boldsymbol{\theta}^{0})$, the score vector mean $\mathbb{E}_{\theta}[\mathbf{s}(\boldsymbol{\theta}^{0};\mathbf{X},Y)|\mathbf{X}]=\int\mathbf{s}(\boldsymbol{\theta}^{0};\mathbf{X},Y=y)\mathbb{P}(Y=y|\mathbf{X};\boldsymbol{\theta})\mathrm{d}y$ will differ from zero. In light of this, the approach of [20] converts concept drift monitoring to the equivalent problem of monitoring for a change in the mean of the score vector $\mathbf{s}(\hat{\boldsymbol{\theta}};\mathbf{X},Y)$, where $\hat{\boldsymbol{\theta}}$ denotes the MLE of $\boldsymbol{\theta}^{0}$ obtained by fitting the model $g(\boldsymbol{\theta};\mathbf{X})$ to the training data. The MLE is given by
@@
The empirical counterpart (2.3) also provides some justification for why the score-based concept drift monitoring approach is effective even when the structure of $\mathbb{P}(Y=y|\mathbf{X};\boldsymbol{\theta})$ induced by the parametric supervised learning model $g(\boldsymbol{\theta};\mathbf{X})$ does not perfectly match the true structure of $\mathbb{P}(Y|\mathbf{X})$. Suppose the predictive relationship between $Y$ and $\mathbf{X}$ for a new set of data $\{(\mathbf{x}_{n+i},y_{n+i}):i=1,2,\cdots,m\}$ differs substantially from what it was over the training data. The value of $\boldsymbol{\theta}$ that minimizes the log-likelihood over the new data will generally differ from the MLE $\hat{\boldsymbol{\theta}}$ over the training data, in which case
@@
$\nabla_{\boldsymbol{\theta}}\frac{1}{m}\sum_{i=1}^{m}\log\mathbb{P}\left(y_{n+i}\middle|\mathbf{x}_{n+i};\boldsymbol{\theta}\right)|_{\boldsymbol{\theta}=\hat{\boldsymbol{\theta}}}=\frac{1}{m}\sum_{i=1}^{m}\mathbf{s}\left(\hat{\boldsymbol{\theta}};\mathbf{x}_{n+i},y_{n+i}\right).$ (2.4)
will differ from zero. The more the predictive relationship for the new data changes relative to the training data, the more we would expect the mean of the new score vectors $\{\mathbf{s}(\hat{\boldsymbol{\theta}};\mathbf{x}_{n+i},y_{n+i}):i=1,2,\cdots,m\}$ evaluated at the same training $\hat{\boldsymbol{\theta}}$ to differ from zero. The same arguments hold if one uses a regularized version of MLE with the score vectors replaced by the derivative of the regularized log-likelihood, which is common in practice: When the predictive relationship changes from the training data, the derivative of the regularized log-likelihood for the new data will generally differ from zero when evaluated at the parameter values that minimize the regularized training log-likelihood.
@@
This was the basis for the concept drift monitoring approach of [20], who used a standard multivariate EWMA (MEWMA) to monitor for changes in the mean of the score vector as the new data are collected. To determine the CL for the MEWMA, they divide the training data $\mathcal{D}=\{(\mathbf{x}_{i},y_{i}):i=1,2,\cdots,n\}$ into two subsets: $\mathcal{D}_{1}:=\{(\mathbf{x}_{i},y_{i}):i=1,2,\cdots,n_{1}\}$ and $\mathcal{D}_{2}:=\{(\mathbf{x}_{i},y_{i}):i=n_{1}+1,n_{1}+2,\cdots,n\}$. The first subset, $\mathcal{D}_{1}$, is used to fit the supervised learning model $g(\hat{\boldsymbol{\theta}};\mathbf{x})$, producing the MLE $\hat{\boldsymbol{\theta}}$. They then compute the score vectors $\mathbf{s}(\hat{\boldsymbol{\theta}};\mathbf{x}_{i},y_{i})$ over the second subset, $\mathcal{D}_{2}$, and apply an MEWMA to these score vectors to empirically compute the CL. Specifically, the CL is taken to be the $1-\alpha$ ($\alpha$ is the desired false-alarm rate) sample quantile of the Hotelling $T^{2}$ statistics, $\{(\mathbf{z}_{i}-\overline{\mathbf{s}})^{\top}\widehat{\boldsymbol{\Sigma}}^{-1}(\mathbf{z}_{i}-\overline{\mathbf{s}}):i=n_{1}+1,n_{1}+2,\cdots,n\}$, where $\overline{\mathbf{s}}=\sum_{i=n_{1}+1}^{n}\mathbf{s}(\hat{\boldsymbol{\theta}};\mathbf{x}_{i},y_{i})/(n-n_{1})$ and $\widehat{\boldsymbol{\Sigma}}=\sum_{i=1}^{n_{1}}(\mathbf{s}(\hat{\boldsymbol{\theta}};\mathbf{x}_{i},y_{i})-\overline{\mathbf{s}})(\mathbf{s}(\hat{\boldsymbol{\theta}};\mathbf{x}_{i},y_{i})-\overline{\mathbf{s}})^{\top}/n_{1}$ are the mean vector and covariance matrix of the score vectors for the second subset, and the MEWMA $\mathbf{z}_{i}$ is defined recursively for $i=n_{1}+1,n_{1}+2,\cdots$ via $\mathbf{z}_{i}=\lambda\mathbf{s}_{i}+(1-\lambda)\mathbf{z}_{i-1}$.
@@
The method of [20] was primarily intended for situations in which $n$ is very large, since the size of $\mathcal{D}_{2}$ must be quite large to accurately determine the $T^{2}$ CL using the above procedure. Our nested bootstrap procedure (described in Section 3) to compute the $T^{2}$ CL results in much more accurate false-alarm rate control. This is especially true when $n$ is not sufficiently large and/or the desired false-alarm rate is small, because the method of [20] requires a very large $\mathcal{D}_{2}$ for small false-alarm rates (e.g., for a desired false-alarm rate of 0.001, the size of $\mathcal{D}_{2}$ must be much larger than 1,000). Moreover, our procedure allows the CL to be computed from the same data to which the supervised learning model is fit. This removes the need to divide the training sample into two subsets and allows the entire sample $\mathcal{D}$ to be used for model fitting, which results in a more accurate model.

### v2

@@ Abstract
Monitoring for changes in a predictive relationship represented by a fitted supervised learning model (i.e., concept drift detection) is a widespread problem in modern data-driven applications. A general and powerful Fisher score–based concept drift approach was recently proposed by Zhang et al. (2023), in which detecting concept drift reduces to detecting changes in the mean of the model’s score vector using a multivariate exponentially weighted moving average (MEWMA). To implement the approach, the initial data must be split into two subsets. The first subset serves as the training sample to which the model is fit, and the second subset serves as an out-of-sample test set from which the MEWMA control limit (CL) is determined. In this paper, we retain the same score-based MEWMA monitoring statistic as Zhang et al. (2023) and focus instead on improving the computation of the control limit. We develop a novel nested bootstrap procedure for calibrating the CL that allows the entire initial sample to be used for model fitting, thereby yielding a more accurate baseline model while eliminating the need for a large holdout set. We show that a standard nested bootstrap substantially underestimates the variability of the monitoring statistic and develop a 0.632-like correction that appropriately accounts for this. We demonstrate the advantages with numerical examples.
Keywords: Control charts, Concept drift, Bootstrap, Predictive modeling, Machine learning
@@ 1 Introduction
The increasing reliance on data-driven decision making has led to the widespread adoption of supervised learning models across various domains. These models aim to capture the predictive relationship $\mathbb{P}(Y|\mathbf{X})$ between a response variable $Y$ and covariates $\mathbf{X}$. However, a fundamental challenge arises when the predictive relationship in new data deviates from that used to train the model, potentially rendering the model’s predictions unreliable or obsolete (Webb et al. 2016, Zhang et al. 2023, Malinovskaya et al. 2024), and/or reflecting a change in process behavior that should be detected. This challenge is particularly acute in domains like finance (Sun et al. 2017) and healthcare analytics (Razak et al. 2023), as most AI systems and algorithms require training data that may contain inherent biases or may not remain representative of the broader population over time (Barocas et al. 2023).
@@
Statistical process monitoring (SPM) and control charts have long served as foundational tools for detecting changes in process characteristics over time (Montgomery 2020). The field has evolved considerably, with recent advances incorporating machine learning methods like kernel methods (Apsemidis et al. 2020) and neural networks (Psarakis 2011) to enhance monitoring capabilities. As datasets have grown larger and organizations increasingly rely on predictive models, the predictive relationship between variables has emerged as a critical characteristic that requires monitoring for quality control purposes, alongside traditional process variables monitoring.
@@
This evolution in the goals of monitoring reflects a fundamental shift in how data are used in practice and/or represents a change in process behavior that should be detected. While traditional SPM methods are designed to detect shifts in the distribution of process variables, modern applications often require understanding changes in the predictive relationships themselves. This phenomenon, known as concept drift in the machine learning literature, occurs when the relationship between input features $\mathbf{X}$ and the target variable $Y$ evolves over time (Webb et al. 2016), potentially degrading model performance (Moreno-Torres et al. 2012, Webb et al. 2016, Žliobaitė et al. 2016) and/or representing a change in process behavior that should be detected. These changes can manifest gradually or abruptly, and may not affect predictive accuracy immediately (Gama et al. 2014), making them difficult to detect using conventional methods.
@@
Most existing concept drift detection methods fall into two categories, with each having significant limitations. The first category consists of error-based approaches that rely primarily on monitoring classification error rates or prediction accuracy metrics (Baena-García et al. 2006, Ross et al. 2012). While straightforward to implement, these methods often fail to detect concept drift when changes in $\mathbb{P}(Y|\mathbf{X})$ do not manifest as increased error rates, such as when decision boundaries shift in ways that maintain similar overall accuracy despite fundamental changes in the underlying relationship. The second category comprises adaptive learning algorithms, which continually retrain or update models to adapt to incoming data streams (Wang et al. 2003, Gama et al. 2014, Krawczyk et al. 2017). While these approaches can improve responsiveness, the adaptive model retraining also requires significant computation and may overfit transient changes.
@@
Related to, but distinctly different from, concept drift detection, profile monitoring methods in statistical process control have been extensively developed to monitor functional relationships between response and predictor variables (Woodall et al. 2004). These methods typically involve monitoring regression parameters or fitted curves over time to detect changes in the relationship between a response variable and one or more explanatory variables (Chang and Yadama 2010, Abbasi et al. 2022). A profile in this context refers to the fitted functional relationship (e.g., a regression curve or model) that characterizes how $Y$ depends on $\mathbf{X}$ within a given batch of data. While this may appear similar to concept drift detection, profile monitoring fundamentally differs in its objectives, its approaches, and the structure of the data to which it applies. In profile monitoring, data are grouped into batches of $(\mathbf{X},Y)$ observations, where some feature of $\mathbb{P}(Y|\mathbf{X})$ (typically $\mathbb{E}[Y|\mathbf{X}]$) for each batch represents a profile associated with the batch, and the objective is to monitor for changes in the nature of the profiles from batch to batch. This requires fitting predictive models separately to each batch of data. In contrast, in concept drift detection, data consist of individual, sequentially arriving $(\mathbf{X},Y)$ observations, and the objective is to detect whether the predictive relationship for new observations changes relative to what it was when a baseline model was fit to prior training data. Changes in the predictive relationship are detected by comparing a new stream of individual $(\mathbf{X},Y)$ observations to the baseline model, which does not involve fitting separate models to the new data.
@@
Recently, Zhang et al. (2023) introduced a new Fisher score based concept drift detection that uses well-established statistical theory to show that detecting changes in $\mathbb{P}(Y|\mathbf{X})$ is equivalent to detecting changes in the mean of the score vector (the gradient of the log-likelihood) of the supervised learning model, a more familiar problem in SPM for which a conventional multivariate exponentially weighted moving average (MEWMA) control charts can be used. Their approach demonstrates superior detection power compared to traditional error-based concept drift methods, while also providing valuable diagnostic information to help understand the nature of detected changes. It also avoids having to continuously retrain the model, as in adaptive learning algorithms. Computing the MEWMA control limit (CL) in Zhang et al. (2023) requires that a second large sample of ($\mathbf{X}$, $Y$) observations be collected, in addition to the training sample to which the baseline model is fit. This can be achieved by splitting the training sample into two sets (with the baseline model fit to the first set, and the second set used to compute the CL), but this reduces the size of the data to which the baseline model is fit and generally requires very large sample sizes to accurately control false-alarm rate. This approach also shares conceptual similarities with risk-adjusted monitoring in healthcare (Steiner et al. 2000), which monitors residuals from a risk model (typically a GLM). However, by using Fisher score vectors, the framework generalizes to any differentiable supervised learning model (including deep neural networks) without requiring the specific parametric forms used in risk-adjusted charts.
@@
Our main contribution is that we present a more data-efficient and reliable approach that uses bootstrapping to compute the CL from the same training sample to which the baseline model is fit, thereby allowing the entire sample to be used to train a more accurate baseline model, as well as providing much more accurate control of false-alarm rate. For this we develop a nested bootstrap procedure with the inner loop accounting for future data variability and outer loop accounting for training sample variability. This is challenging because a naïve nested bootstrap procedure substantially underestimates variability of the MEWMA monitoring statistic for reasons that we discuss later. To account for this, we develop a 0.632-like (analogous to, but different from the original 0.632 correction of Efron (1983)) that controls false-alarm rate much more accurately than the two-sample CL calculation of Zhang et al. (2023), especially when the training sample size and/or the desired false-alarm rate is small. Our implementation takes advantage of the inherent parallelism in the nested bootstrap structure, making the method computationally efficient and practical for modern applications involving complex models like deep neural networks.
We note that the detection statistic utilized in our approach is identical to the score-based MEWMA statistic proposed by Zhang et al. (2023). Extensive empirical comparisons in Zhang et al. (2023) have already demonstrated that this statistic yields superior detection power compared to error-based methods (e.g., DDM (Gama et al. 2004), EDDM (Baena-García et al. 2006)) and adaptive windowing, particularly for drifts that do not immediately degrade prediction accuracy. Therefore, our numerical analysis focuses specifically on the accuracy of the proposed bootstrap Control Limits (Type I error control) and data efficiency, rather than repeating the detection power comparisons established in prior literature.
@@ 2 Background on Score-Based Concept Drift Detection
This work assumes a parametric supervised learning $g(\boldsymbol{\theta};\mathbf{X})$ is used to represent the conditional distribution $\mathbb{P}(Y|\mathbf{X};\boldsymbol{\theta})$ of the target variable $Y$ given the input features $\mathbf{X}\in\mathbb{R}^{p}$, where $\boldsymbol{\theta}$ denotes the parameters of the supervised learning, and $p$ is the input dimension. For example, for classification problems, $g(\boldsymbol{\theta};\mathbf{X})$ directly outputs the response class probabilities; and for regression with Gaussian errors and squared-error loss, $g(\boldsymbol{\theta};\mathbf{X})$ represents the conditional mean $\mathbb{E}[Y|\mathbf{X};\boldsymbol{\theta}]$ of a Gaussian $\mathbb{P}(Y|\mathbf{X};\boldsymbol{\theta})$. As in Zhang et al. (2023), we assume $g(\boldsymbol{\theta};\mathbf{X})$ is fit to the training data via maximum likelihood estimation (MLE), perhaps with regularization. This parametric assumption implies that the method is applicable to models such as generalized linear models (GLMs) and neural networks (where gradients with respect to weights are well-defined). It is not directly applicable to non-parametric or tree-based methods like Random Forests or XGBoost, which do not admit a differentiable likelihood function with respect to fixed parameters.
Denote the training data to which $g(\boldsymbol{\theta};\mathbf{X})$ is fit by $\{(\mathbf{x}_{i},y_{i}):i=1,2,\cdots,n\}$, assumed to be an i.i.d. sample drawn from the joint distribution $\mathbb{P}(Y,\mathbf{X})$. For each observation $(\mathbf{x}_{i},y_{i})$, the (Fisher) score vector is defined as $\mathbf{s}(\boldsymbol{\theta};\mathbf{x}_{i},y_{i})=\nabla_{\boldsymbol{\theta}}\log\mathbb{P}(y_{i}|\mathbf{x}_{i};\boldsymbol{\theta})$. Under certain regularity conditions, if the assumed parametric model is indeed the true data-generating mechanism with true parameters denoted by $\boldsymbol{\theta}^{0}$, a fundamental result in statistical theory (Bickel and Doksum 2015)[Proposition 3.4.4] states that the expected value of the score vector, when evaluated at $\boldsymbol{\theta}^{0}$, is equal to zero
@@
In the context of machine learning, the notion of concept drift refers to the phenomenon where the underlying relationship between $\mathbf{X}$ and $Y$ evolves over time, which can be characterized as a shift in $\mathbb{P}(Y|\mathbf{X};\boldsymbol{\theta})$. In the parametric setting of Zhang et al. (2023), concept drift translates to a change in $\boldsymbol{\theta}$. Under certain identifiability conditions, when the parameters change (to some $\boldsymbol{\theta}\neq\boldsymbol{\theta}^{0})$, the score vector mean $\mathbb{E}_{\theta}[\mathbf{s}(\boldsymbol{\theta}^{0};\mathbf{X},Y)|\mathbf{X}]=\int\mathbf{s}(\boldsymbol{\theta}^{0};\mathbf{X},Y=y)\mathbb{P}(Y=y|\mathbf{X};\boldsymbol{\theta})\mathrm{d}y$ will differ from zero. In light of this, the approach of Zhang et al. (2023) converts concept drift monitoring to the equivalent problem of monitoring for a change in the mean of the score vector $\mathbf{s}(\hat{\boldsymbol{\theta}};\mathbf{X},Y)$, where $\hat{\boldsymbol{\theta}}$ denotes the MLE of $\boldsymbol{\theta}^{0}$ obtained by fitting the model $g(\boldsymbol{\theta};\mathbf{X})$ to the training data. The MLE is given by
@@
The empirical counterpart (2.3) also provides some intuition for why the score-based concept drift monitoring approach can remain effective even when the parametric supervised learning model $g(\boldsymbol{\theta};\mathbf{X})$ is only an approximation to the true predictive relationship $\mathbb{P}(Y\mid\mathbf{X})$. Suppose the predictive relationship between $Y$ and $\mathbf{X}$ for a new set of data $\{(\mathbf{x}_{n+i},y_{n+i}):i=1,2,\cdots,m\}$ differs substantially from what it was over the training data. The value of $\boldsymbol{\theta}$ that minimizes the log-likelihood over the new data will generally differ from the MLE $\hat{\boldsymbol{\theta}}$ over the training data, in which case
@@
$\nabla_{\boldsymbol{\theta}}\frac{1}{m}\sum_{i=1}^{m}\log\mathbb{P}\left(y_{n+i}\middle|\mathbf{x}_{n+i};\boldsymbol{\theta}\right)|_{\boldsymbol{\theta}=\hat{\boldsymbol{\theta}}}=\frac{1}{m}\sum_{i=1}^{m}\mathbf{s}\left(\hat{\boldsymbol{\theta}};\mathbf{x}_{n+i},y_{n+i}\right),$ (2.4)
will differ from zero. This argument does not require the supervised learning model to be correctly specified. Even under model misspecification, the maximum likelihood estimator converges to a pseudo-true parameter minimizing the Kullback–Leibler divergence. When the underlying predictive relationship changes, this pseudo-true parameter typically shifts, implying that the expected score vector evaluated at the original training fit is no longer zero. The same arguments hold if one uses a regularized version of MLE with the score vectors replaced by the derivative of the regularized log-likelihood, which is common in practice. This robustness property is supported by the empirical results in Zhang et al. (2023) and subsequent follow-ups, including applications to microstructure image data in Zhang et al. (2021) (for which the supervised learning model is always only an approximate representation of the true relationship). It is also demonstrated in our nonlinear oscillator example (Section 4.2), where a simplified neural network only approximates the complex physical dynamics but still allows changes in the physical parameters to be successfully detected. Theorem 3.4 formalizes this detectability argument under model misspecification.
@@
This was the basis for the concept drift monitoring approach of Zhang et al. (2023), who used a standard multivariate EWMA (MEWMA) to monitor for changes in the mean of the score vector as the new data are collected. To determine the CL for the MEWMA, they divide the training data $\mathcal{D}=\{(\mathbf{x}_{i},y_{i}):i=1,2,\cdots,n\}$ into two subsets: $\mathcal{D}_{1}:=\{(\mathbf{x}_{i},y_{i}):i=1,2,\cdots,n_{1}\}$ and $\mathcal{D}_{2}:=\{(\mathbf{x}_{i},y_{i}):i=n_{1}+1,n_{1}+2,\cdots,n\}$. The first subset, $\mathcal{D}_{1}$, is used to fit the supervised learning model $g(\hat{\boldsymbol{\theta}};\mathbf{x})$, producing the MLE $\hat{\boldsymbol{\theta}}$. They then compute the score vectors $\mathbf{s}(\hat{\boldsymbol{\theta}};\mathbf{x}_{i},y_{i})$ over the second subset, $\mathcal{D}_{2}$, and apply an MEWMA to these score vectors to empirically compute the CL. Specifically, the CL is taken to be the $1-\alpha$ ($\alpha$ is the desired false-alarm rate) sample quantile of the Hotelling $T^{2}$ statistics, $\{(\mathbf{z}_{i}-\overline{\mathbf{s}})^{\top}\widehat{\boldsymbol{\Sigma}}^{-1}(\mathbf{z}_{i}-\overline{\mathbf{s}}):i=n_{1}+1,n_{1}+2,\cdots,n\}$, where $\overline{\mathbf{s}}=\sum_{i=n_{1}+1}^{n}\mathbf{s}(\hat{\boldsymbol{\theta}};\mathbf{x}_{i},y_{i})/(n-n_{1})$ and $\widehat{\boldsymbol{\Sigma}}=\sum_{i=1}^{n_{1}}(\mathbf{s}(\hat{\boldsymbol{\theta}};\mathbf{x}_{i},y_{i})-\overline{\mathbf{s}})(\mathbf{s}(\hat{\boldsymbol{\theta}};\mathbf{x}_{i},y_{i})-\overline{\mathbf{s}})^{\top}/n_{1}$ are the mean vector and covariance matrix of the score vectors for the second subset, and the MEWMA $\mathbf{z}_{i}$ is defined recursively for $i=n_{1}+1,n_{1}+2,\cdots$ via $\mathbf{z}_{i}=\lambda\mathbf{s}_{i}+(1-\lambda)\mathbf{z}_{i-1}$, where $\lambda$ is the MEWMA smoothing parameter.
@@
The method of Zhang et al. (2023) was primarily intended for situations in which $n$ is very large, since the size of $\mathcal{D}_{2}$ must be quite large to accurately determine the $T^{2}$ CL using the above procedure. Our nested bootstrap procedure (described in Section 3) to compute the $T^{2}$ CL results in much more accurate false-alarm rate control. This is especially true when $n$ is not sufficiently large and/or the desired false-alarm rate is small, because the method of Zhang et al. (2023) requires a very large $\mathcal{D}_{2}$ for small false-alarm rates (FAR) (e.g., for a desired false-alarm rate of 0.001, the size of $\mathcal{D}_{2}$ must be much larger than 1,000). Moreover, our procedure allows the CL to be computed from the same data to which the supervised learning model is fit. This removes the need to divide the training sample into two subsets and allows the entire sample $\mathcal{D}$ to be used for model fitting, which results in a more accurate model.
Practitioners often encounter scenarios where concept drift affects the relationship between $Y$ and only a subset of the feature vector $\mathbf{X}$ (e.g., the example in Section 4.2, later). In such cases, the true parameters associated with the unaffected features may remain constant, while those associated with the drifting features change. Because the score vector $\mathbf{s}(\hat{\boldsymbol{\theta}};\mathbf{x},y)$ contains components for all model parameters, and the MEWMA statistic $T^{2}$ aggregates deviations across the entire parameter vector space, the method can detect these partial shifts. The detection power will of course depend on how strongly the drifting subset influences the overall likelihood.